• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Lindgren

    Log entry time 18:37:15 on April 5,2008

    Entry number 230129

    keyword=Shift Summary

    Day Shift Summary
    Start 8:09 Run 2325 HRS=20.5, C Target, BB=390 A Events = 0.220M
    Start 8:27 Run 2326-2327 HRS=20.5, LH2 Target, BB=390 A pi0 Test
    Start 8:38 Runs 2328-2329 HRS=20.5, LH2 Target, Dead Time Test
    good data below
    Start 8:45 Runs 2330 HRS=20.5, LH2 Target, pi0 Events=0.48M
    Start 8:45 Runs 2331 HRS=20.5, LH2 Target, pi0 Events=0.42M

    Control Access
    See comments in Halog by Brad and others
    Enter Hall to change gas on drift chambers and lower BigBite Sieve Slit. Note Balloons are still inflated.
    Back to Beam permit
    Move HRS from 20.5 to 35.5
    No beam until about 1400.
    Beam back about 1400
    Put in C holey target, Check raster with spot++ Beam centered very well.

    Take elastic data in HRS and Recoil protons in BigBite
    With SIEVE IN
    Start 14:26 Run 2340 HRS=35.5, LH2 Target, BB=390 A Events = 0.5M
    Start 15:26 Run 2341 HRS=35.5, LH2 Target, BB=390 A Events = 0.5M
    (Event Recorder Crashed)
    Start 15:49 Run 2342 HRS=35.5, LD2 Target, BB=390 A Events = Junk
    (ROC5) Crashed
    Run 2343 - 2344 ROC5 Crashed
    Good data below
    Start 16:21 Run 2345 HRS=35.5, LD2 Target, BB=390 A Events = 0.43M
    Start 16:27 Run 2346 HRS=35.5, LD2 Target, BB=390 A Events = 0.52M